# Prediction of Energy Resolution in the JUNO Experiment

### v1

@@ 8 Summary
The energy resolution is a crucial parameter in determining the sensitivity of the JUNO experiment to NMO. However, in addition to the statistical fluctuations in the detected number of PE, several other effects impact the energy resolution in the detection of the IBD signals. This paper presents a comprehensive study of the energy resolution in the JUNO experiment, incorporating the latest knowledge and updates in the detector construction stage. This includes a better understanding of the detector structures, more precise measurements of the optical properties of the LS, comprehensive evaluations of the characteristics of both LPMTs and SPMTs, improved modeling of the spectral and angular dependencies of the LPMTs’ PDE, and better constraints on the absolute scintillation and Cherenkov light yield based on data from the Daya Bay experiment. All of these updates have been implemented in the JUNO detector simulation, which also includes detailed modeling of the PMT responses and readout electronics in the electronics simulation.
By using data samples generated with the full detector and electronics simulation, a full-chain data processing of calibration and reconstruction is performed to evaluate factors that can smear the energy resolution. These factors include residual energy non-uniformity after reconstruction, accuracy of reconstructed energies from reconstruction algorithms, impacts of dark count rate and SPE charge resolution, among others. After considering these factors, it is found that an overall energy resolution of 2.95% at 1 MeV can be achieved for positrons from IBD signals, and the obtained energy resolution curve has been applied in the recent NMO analysis of JUNO. Furthermore, the contribution of major effects in the energy resolution budget is estimated. This study serves as a reference for interpreting future measurements of energy resolution in JUNO data taking and provides a guideline for understanding the energy resolution of LS-based detectors.
After the data-taking phase of the JUNO experiment begins, further updates and improvements are expected in the understanding of the energy resolution. The analysis of the collected data will provide valuable insights and allow for refinements in the modeling and calibration of the detector. These updates and improvements will contribute to the ongoing efforts to optimize the performance of the JUNO detector and enhance its sensitivity to NMO.
@@ Acknowledgement
We acknowledge the Daya Bay collaboration for providing the Monte Carlo simulation software. We are grateful for the ongoing cooperation from the China General Nuclear Power Group. This work was supported by the Chinese Academy of Sciences, the National Key R&D Program of China, the CAS Center for Excellence in Particle Physics, Wuyi University, and the Tsung-Dao Lee Institute of Shanghai Jiao Tong University in China, the Institut National de Physique Nucléaire et de Physique de Particules (IN2P3) in France, the Istituto Nazionale di Fisica Nucleare (INFN) in Italy, the Italian-Chinese collaborative research program MAECI-NSFC, the Fond de la Recherche Scientifique (F.R.S-FNRS) and FWO under the “Excellence of Science – EOS” in Belgium, the Conselho Nacional de Desenvolvimento Científico e Tecnològico in Brazil, the Agencia Nacional de Investigacion y Desarrollo in Chile, the Charles University Research Centre and the Ministry of Education, Youth, and Sports in Czech Republic, the Deutsche Forschungsgemeinschaft (DFG), the Helmholtz Association, and the Cluster of Excellence PRISMA+ in Germany, the Joint Institute of Nuclear Research (JINR) and Lomonosov Moscow State University in Russia, the joint Russian Science Foundation (RSF) and National Natural Science Foundation of China (NSFC) research program, the MOST and MOE in Taiwan, the Chulalongkorn University and Suranaree University of Technology in Thailand, and the University of California at Irvine in USA.

### v2

@@ 8 Summary
The energy resolution is a crucial parameter in determining the sensitivity of the JUNO experiment to NMO. However, in addition to the statistical fluctuations in the detected number of PEs, several other effects impact the energy resolution in the detection of IBD signals. This paper presented a comprehensive study of the energy resolution in the JUNO experiment, incorporating the latest knowledge and updates in the detector construction stage. This includes a better understanding of the detector structures, more precise measurements of the optical properties of the LS, comprehensive evaluations of the characteristics of both LPMTs and SPMTs, improved modeling of the spectral and angular dependencies of the LPMTs’ PDE, and better constraints on the absolute scintillation and Cherenkov light yield based on data from the Daya Bay experiment. All of these updates have been implemented in the JUNO detector simulation, which also includes detailed modeling of the PMT responses and readout electronics in the electronics simulation.
By using data samples generated with the full detector and electronics simulation, a full-chain data processing of calibration and reconstruction was performed to evaluate factors that can smear the energy resolution. These factors include residual energy non-uniformity after reconstruction, accuracy of reconstructed energies from reconstruction algorithms, impacts of dark count rate, and SPE charge resolution. After considering these factors, it was found that an overall energy resolution of 2.95% at 1 MeV can be achieved for positrons from IBD signals, and the obtained energy resolution curve has been applied in the recent NMO analysis of JUNO. Furthermore, the contribution of major effects in the energy resolution budget was estimated. This study serves as a reference for interpreting future measurements of energy resolution in JUNO data collection and provides a guideline for understanding the energy resolution of LS-based detectors.
After the data collection phase of the JUNO experiment begins, further updates and improvements are expected in the understanding of the energy resolution. The analysis of the collected data will provide valuable insights and allow for refinements in the modeling and calibration of the detector. These updates and improvements will contribute to the ongoing efforts to optimize the performance of the JUNO detector and enhance its sensitivity to NMO.
@@ Acknowledgement
We acknowledge the Daya Bay collaboration for providing the Monte Carlo simulation software. We are grateful for the ongoing cooperation from the China General Nuclear Power Group. This work was supported by the Chinese Academy of Sciences, the National Key R&D Program of China, the CAS Center for Excellence in Particle Physics, Wuyi University, and the Tsung-Dao Lee Institute of Shanghai Jiao Tong University in China, the Institut National de Physique Nucléaire et de Physique de Particules (IN2P3) in France, the Istituto Nazionale di Fisica Nucleare (INFN) in Italy, the Italian-Chinese collaborative research program MAECI-NSFC, the Fond de la Recherche Scientifique (F.R.S-FNRS) and FWO under the “Excellence of Science – EOS” in Belgium, the Conselho Nacional de Desenvolvimento Científico e Tecnològico in Brazil, the Agencia Nacional de Investigacion y Desarrollo and ANID Millennium Science Initiative Program — ICN2019_044 in Chile, the Charles University Research Centre and the Ministry of Education, Youth, and Sports in Czech Republic, the Deutsche Forschungsgemeinschaft (DFG), the Helmholtz Association, and the Cluster of Excellence PRISMA+ in Germany, the Joint Institute of Nuclear Research (JINR) and Lomonosov Moscow State University in Russia, the joint Russian Science Foundation (RSF) and National Natural Science Foundation of China (NSFC) research program, the MOST and MOE in Taiwan, the Chulalongkorn University and Suranaree University of Technology in Thailand, and the University of California at Irvine and the National Science Foundation in USA.